You want to save an equal amount each year for the next 38 years, at which time you will retire. What amount of annual savings are needed if you desire a retirement income of $55,000 a year for 25 years and earn 7.5 percent, compounded annually? Select one:

A. $2,889.04

B. $3,333.33

C. $2,640.85

D. $3,406.16

E. $3,146.32
Answer #1

Answer: Option E: $3146.32

Total of PV of Retirement income at end of 38 years= 55000*(1-(1+7.5%)^-25)/7.5%=$613082.02

Let be X saving each year

So FV of X each year =613082.02

X*((1+7.5%)^38-1)/7.5%=613082.02

X=$3146.32